# Quantum Simulation of Ligand-like Molecules through Sample-based Quantum Diagonalization in Density Matrix Embedding Framework

### v2

@@ Abstract
The accurate treatment of electron correlation in extended molecular systems remains computationally challenging using classical electronic structure methods. Hybrid quantum–classical algorithms offer a potential route to overcome these limitations; however, their practical deployment on existing quantum computers requires strategies that both reduce problem size and mitigate hardware noise. In this work, we combine Density Matrix Embedding Theory (DMET) with Sample-based Quantum Diagonalization (SQD) to compute ground-state energies of a set of natural ligand-like molecules in the minimal Slater Type Orbital (STO-3G) basis set. DMET provides a systematic fragmentation of a molecule into embedded impurity subproblems, while SQD enables construction and classical diagonalization of reduced configuration spaces through quantum sampling enhanced by iterative configuration recovery. The resulting embedded Hamiltonians are solved on IBM’s Eagle R3 superconducting quantum hardware (IBM Sherbrooke). The DMET-SQD energies obtained for all systems considered exhibit strong agreement with DMET-FCI benchmark values within chemical accuracy (1 kcal/mol). These results demonstrate that sample-based quantum methods, when integrated with a robust embedding framework, can reliably extend quantum computation towards simulation of chemically relevant molecular systems, showcasing potential applications in the field of drug discovery.
@@ 1 Introduction
Classical fragmentation techniques offer a practical route to handling the exponential scaling of electronic structure methods by dividing large molecular systems into smaller subsystems. This fragment-based viewpoint allows chemically realistic systems to be studied without treating the entire system at once, while still retaining a meaningful description of inter-fragment interactions. By decomposing the global wavefunction or Hamiltonian into fragment-level contributions, these approaches enable the treatment of chemically realistic systems while retaining a controlled description of inter-fragment interactions. Prominent examples include Density Matrix Embedding Theory (DMET) [95, 51, 69, 22] and its multi-reference variants [77, 92], Divide and Conquer (DC) method [68, 6], and the Fragment Molecular Orbital (FMO) framework [30], all of which have demonstrated practical applicability across a wide range of molecular and materials problems [97]. These techniques provide a systematic route for reducing the effective dimensionality of quantum chemical calculations by localizing electronic structure in real space or orbital subspaces, while incorporating environmental effects through embedding potentials, density constraints, or effective Hamiltonians.
@@
Beyond these foundational approaches, several modern fragmentation schemes aim to preserve higher-order correlation effects while maintaining computational efficiency [13, 14, 40, 72, 70, 60, 96]. Coupled-cluster downfolding methods construct effective low-dimensional Hamiltonians by integrating out inactive degrees of freedom using classical CC theory, yielding size-consistent and systematically improvable reduced models [13, 14, 52]. Variational Localized Active Space methods, such as LASSCF [40, 72], explicitly partition the active space into weakly entangled local subspaces that can be optimized independently, providing a natural interface for hybrid classical-quantum workflows. Related embedding paradigms include wavefunction-based embedding methods (EWF) [70], Many-Body Expansion (MBE) [60, 96], and Dynamical Mean-Field Theory (DMFT) [35, 17], which employ different strategies for capturing non-local correlation and dynamical effects. More recently, Bootstrap Embedding (BE) [59] has been introduced as a hybrid quantum-classical embedding framework that employs a self-consistent feedback loop between quantum impurity solvers and an effective classical environment.
@@
In parallel, several approaches have focused on adapting classical fragmentation techniques for use with quantum simulation algorithms. Notable examples include DMET-VQE implementations on superconducting and trapped-ion platforms [49, 9], as well as DMET-SQD [86] and EWF-SQD [85] approaches demonstrated on IBM quantum hardware. Related efforts such as divide-and-conquer VQE (DC-VQE) [33] and fragment molecular orbital extensions to quantum solvers, including Effective FMO-VQE [57], further illustrate the breadth of strategies combining classical fragmentation with quantum sampling or variational techniques. Collectively, these works define a growing ecosystem of hybrid methods that aim to balance accuracy, scalability, and near-term hardware constraints in quantum-enabled electronic structure calculations.
SQD has gained significant interest in recent times [82, 75, 74, 98] and is subject to further exploration to develop and assess its performance within fragmentation frameworks [86, 85, 16]. Towards this, we develop the sample-based technique, SQD within the DMET framework and systematically study its ability in accurate prediction of ground-state energetics across a series of ligand-like molecules. These ligand-like molecules exhibit reduced symmetry in contrast to the highly symmetric model systems considered in previous studies [86]. As a result, electron delocalization over the entire molecule introduces qualitatively different challenges for embedding and sampling, providing a rigorous test of the robustness and practical applicability of the proposed approach. The ligand-like molecules considered for the study include: Cyanic Acid($HOCN$), Formaldehyde Oxime ($CH_{3}NO$), O-methyl-hydroxylamine ($CH_{5}NO$), Methyl Isocyanate ($C_{2}H_{3}NO$), Acetaldehyde Oxime ($C_{2}H_{5}NO$), Carbamide/Urea ($CH_{4}N_{2}O$), Nitrosyl Chloride ($NOCl$), and Hydroxythiocyanate ($HOSCN$). All of these compounds have molecular weights below 76 Da (see Fig. 4.1) and together span a chemically diverse set of small, pharmacologically motivated molecular systems. Several members of this set, particularly urea derivatives and methoxyamine-related systems have established relevance in medicinal chemistry [20, 26, 36, 58]. This makes the collection an appropriate testbed for assessing sample-based quantum simulation techniques in the context of drug-discovery-oriented molecular systems.
@@
Concisely detailing the work, in Section 2 an overview of the methods used are described briefly, in Section 3 the experimental details and methodology are discussed, in Section 4 the results obtained are discussed and finally in Section 5 we conclude the work and mention future prospects.
@@ 2.2 Density Matrix Embedding Theory
Figure 2.1: Illustrative example of bath-orbital construction for the $H_{2}O$ molecule in the STO-3G basis. The system is fragmented so that the orbital localized on a single hydrogen atom constitutes the fragment. (a) shows an iso-surface of this fragment orbital; (b) shows the corresponding bath orbital obtained via DMET, which is a linear combination of the remaining localized spatial orbitals in the molecular orbital basis. Together, (a) and (b) span the impurity subspace for the H-fragment.
@@ 2.3 Sample-based Quantum Diagonalization
SQD [82] is a combination of the quantum sampling method QSCI [47] and an error-mitigation technique introduced in the work, S-CoRe, which “recovers” the noisy configurations. These configurations thus obtained create a sub-space, which is further expanded, and then the projection sub-space is obtained over which SCI [28] is performed to obtain the ground state energy.
@@ 3.1 DMET-SQD Workflow
    The molecular co-efficient matrix $C$, and the HF determinant (in the case of single-reference systems) of the whole system are obtained using Restricted HF method (here we only encounter molecular systems with fully filled spatial orbitals).
@@
    The bath orbitals are constructed with the remaining environment orbitals for the fragment $A_{y}$. Here, the occupation threshold $\epsilon_{occ}$ is set to $10^{-13}$ for discarding virtual orbitals [95].
@@
    The Fock matrix $\hat{F}_{imp}$ (where $imp=A_{y}+B_{y}$) for each of the impurities is obtained, and the chemical potential $\mu_{glob}$ is subtracted from the diagonal entries corresponding to the fragment orbitals:
@@
    Using $\hat{F}_{imp}^{\mathrm{new}}(\mu_{\mathrm{glob}})$ from Eq. (12), an SCF calculation is carried out in the impurity space to obtain updated molecular properties, such as the impurity coefficient matrix $C_{imp}^{\mathrm{new}}(\mu_{\mathrm{glob}})$ and the corresponding one-particle density matrix $D^{y}(\mu_{\mathrm{glob}})$. The standard convention is to initialize $\mu_{glob}^{init}=0$.
@@
    Now, we use SQD, as the high-level solver to estimate the ground state energy and eigen-vector of these "impurities". The one-electron ($h_{imp}^{pq}(\mu_{glob})$) and two-electron ($h_{imp}^{pqrs}(\mu_{glob})$) Hamiltonian terms (where $pqrs\in imp$) obtained after the localization, serve as the $\hat{H}_{emb}(\mu_{glob})$, and the configurations obtained through quantum sampling (and subsequent Configuration Recovery) are used to form the subspace onto which $\hat{H}_{emb}$ is projected. In the case of calculating the reference energies, this step would be replaced by FCI, without any loss of generalization.
@@
    Using the eigen-vector obtained in the previous step, the following quantities are calculated:
@@
    Now similar to the fragment $A_{y}$, the steps 2 to 6 are repeated for all the remaining $N_{frag}-1$ fragments, where $N_{frag}$ represents the number of fragments in the system. The total energy $E_{tot}(\mu_{glob})$, the number of electrons of the system $N_{tot}$, and the error in electron number $N^{ele}_{err}$ can be thus obtained by:
@@
    Next, the root of the function $N_{err}^{ele}(\mu_{glob})$ is computed to within the convergence threshold $\epsilon_{conv}$, by iteratively updating $\mu_{glob}$, and repeating the steps 2 to 7, using a root finding method (in this case we use Newton-Secant method [73]). In other words, we wish to obtain a $\mu_{glob}^{\star}$ such that:
@@ 3.2 Experimental Details
The molecules taken for the study are shown in Sec. B. The STO–3G basis set was employed in this work, due to its minimal size and computational efficiency, which make it particularly well suited for quantum simulations on current NISQ-era hardware, and the fragmentation was performed such that each atom in the fragment is treated as a fragment for all the molecules.
@@
For this reason, FCI is employed as the benchmark solver, as it accesses the full configuration space and is both size-consistent and size-extensive. The minimal basis set and force-field-optimized geometries were chosen to enable a controlled methodological benchmarking of DMET-SQD against DMET-FCI under current NISQ-era hardware and access constraints, where qubit counts, circuit depths, and limited execution time make the use of larger basis sets impractical. This choice further allows FCI to be performed without active-space truncation, ensuring that any degradation in correlation arises from the embedding and solver strategy, and thereby validating the ability of DMET-SQD to reproduce the correlation capture of DMET-FCI within the bounds of the present study.
The number of shots used for quantum sampling was $10^{4}$ and was kept consistent across all the impurities, regardless of the number of spin orbitals within them. For each impurity, 5 iterations of S-CoRe were performed. The maximum allowed samples per batch was set to $10^{5}$. The circuits were transpiled [56] with optimization set to level 3, and no additional error handling techniques were employed in these series of experiments. The convergence threshold $\epsilon_{conv}$ for $\mu_{glob}$ was set to $1.48\times 10^{-8}$, which is the default parameter in scipy.optimize.newton method.
@@ 4 Results and Discussion
The details pertaining to the fragmentation settings, quantum resource requirements, and the dimensions of the symmetry space and Hilbert space are summarized in Table 4.1. The resource estimation is presented in the IBM Sherbrooke superconducting hardware basis gate set {$X$, $R_{z}$, $\sqrt{X}(S_{x})$, ECRECR}. The largest quantum-sampling experiment employed 30 qubits, with a symmetry-space dimension of 25,050,025 and a Hilbert-space dimension of 1,073,741,824. The largest quantum circuit depth was 1081. These entries have been highlighted in the table to indicate the maximal values observed in our study.

### v3

@@ Abstract
The accurate treatment of electron correlation in extended molecular systems remains computationally challenging using classical electronic structure methods. Hybrid quantum-classical algorithms offer a potential route to overcome these limitations; however, their practical deployment on existing quantum computers requires strategies that both reduce problem size and mitigate hardware noise. In this work, we investigate ground-state energy calculations of ligand-like molecules using Sample-based Quantum Diagonalization (SQD) within the Density Matrix Embedding Theory (DMET) framework, focusing on low-symmetry systems with diverse bonding motifs which exhibit subsystem-dependent variations in fragment-environment entanglement. These entanglement-based variations directly influence bath orbital construction, impurity sizes, and the structure of the embedded Hamiltonians, posing nontrivial challenges for both embedding and quantum sampling. By combining DMET fragmentation with SQD-based construction of reduced configuration spaces through quantum sampling and iterative configuration recovery, we perform quantum simulations on IBM’s Eagle R3 superconducting quantum hardware(IBM Sherbrooke) thereby, showing that the entanglement structure across embedding subsystems plays a central role in determining the efficiency and accuracy of the simulations. Despite these complexities, we show that the DMET-SQD framework yields ground-state energies in strong agreement with DMET-FCI benchmarks, achieving chemical accuracy (1 kcal/mol) across all systems studied. These results demonstrate that SQD-based quantum simulations can be robustly extended to low-symmetry, chemically realistic, industry relevant molecules, and highlight the importance of entanglement-aware embedding strategies for scalable quantum electronic structure calculations.
@@ 1 Introduction
Classical fragmentation techniques offer a practical route to handling the exponential scaling of electronic structure methods by dividing large molecular systems into smaller subsystems. This fragment-based viewpoint allows chemically realistic systems to be studied without treating the entire system at once, while still retaining a meaningful description of inter-fragment interactions. By decomposing the global wavefunction or Hamiltonian into fragment-level contributions, these approaches enable the treatment of chemically realistic systems while retaining a controlled description of inter-fragment interactions. Prominent examples include Density Matrix Embedding Theory (DMET) [96, 51, 69, 21] and its multi-reference variants [77, 93], Divide and Conquer (DC) method [68, 6], and the Fragment Molecular Orbital (FMO) framework [29], all of which have demonstrated practical applicability across a wide range of molecular and materials problems [98]. These techniques provide a systematic route for reducing the effective dimensionality of quantum chemical calculations by localizing electronic structure in real space or orbital subspaces, while incorporating environmental effects through different mechanisms: embedding potentials in density-based approaches such as DFT embedding, density matrix matching in DMET, and effective Hamiltonians in wavefunction-based embedding frameworks [89, 96].
@@
Beyond these foundational approaches, several modern fragmentation schemes aim to preserve higher-order correlation effects while maintaining computational efficiency [12, 13, 39, 72, 70, 60, 97]. Coupled-cluster downfolding methods construct effective low-dimensional Hamiltonians by integrating out inactive degrees of freedom using classical CC theory, yielding size-consistent and systematically improvable reduced models [12, 13, 52]. Variational Localized Active Space methods, such as LASSCF [39, 72], explicitly partition the active space into weakly entangled local subspaces that can be optimized independently, providing a natural interface for hybrid classical-quantum workflows. Related embedding paradigms include wavefunction-based embedding methods (EWF) [70], Many-Body Expansion (MBE) [60, 97], and Dynamical Mean-Field Theory (DMFT) [34, 16], which employ different strategies for capturing non-local correlation and dynamical effects. More recently, Bootstrap Embedding (BE) [59] has been introduced as a hybrid quantum-classical embedding framework that builds upon DMET by enforcing consistency of reduced density matrices across fragment boundaries, matching the density matrices of the central region of a fragment with those of its neighboring fragments.
@@
In parallel, several approaches have focused on adapting classical fragmentation techniques for use with quantum simulation algorithms. Notable examples include DMET-VQE implementations on superconducting and trapped-ion platforms [49, 48, 43], as well as DMET-SQD [86] and EWF-SQD [85] approaches demonstrated on IBM quantum hardware. Related efforts such as divide-and-conquer VQE (DC-VQE) [32] and fragment molecular orbital extensions to quantum solvers, including Effective FMO-VQE [57], further illustrate the breadth of strategies combining classical fragmentation with quantum sampling or variational techniques. Collectively, these works define a growing ecosystem of hybrid methods that aim to balance accuracy, scalability, and near-term hardware constraints in quantum-enabled electronic structure calculations.
SQD has gained significant interest in recent times [82, 75, 74, 99] and is subject to further exploration to develop and assess its performance within fragmentation frameworks [86, 85, 15]. Towards this, we assess the feasibility of SQD within the DMET framework for ligand-like molecules extending upon the recently introduced DMET-SQD framework [86] for strongly correlated symmetric systems. Herein, we extend this framework to accomodate for a wider class of molecules which are biologically relevant, unlike the previously studied systems [86] which had high point group symmetry, these ligand-like molecules belong to low-symmetry classes (typically $C_{1}$).. This makes the universal thresholding criteria to pick the dominant configurations over individual embedding problems a challenging task, a difficulty which, provides a rigorous test of the robustness and practical applicability of the proposed approach.
The ligand-like molecules considered for the study include: Cyanic Acid($HOCN$), Formaldehyde Oxime ($CH_{3}NO$), O-methyl-hydroxylamine ($CH_{5}NO$), Methyl Isocyanate ($C_{2}H_{3}NO$), Acetaldehyde Oxime ($C_{2}H_{5}NO$), Carbamide/Urea ($CH_{4}N_{2}O$), Nitrosyl Chloride ($NOCl$), and Hydroxythiocyanate ($HOSCN$). All of these compounds have molecular weights below 76 Da (see Fig. 4.1) and together span a chemically diverse set of small, pharmacologically motivated molecular systems. Several members of this set, particularly urea derivatives and methoxyamine-related systems have established relevance in medicinal chemistry [19, 25, 35, 58]. Further more, from a theoretical point of view, due to the heterogeneous nature of individual embedding problems, unlike the previous studies, it is important to assess the entanglement of the fragment orbitals with the associated bath orbitals. This makes a straight-forward implementation of DMET-SQD a challenging exercise and further refinement of the DMET bath is required, that affects sampling complexity, and this will be further elaborated in Section 4. This makes the chosen set of molecules an appropriate testbed for assessing DMET-SQD in the context of biologically and industry relevant molecular systems.
@@
Concisely detailing the work, in Section 2 an overview of the methods used are described briefly, in Section 3 the experimental details and methodology are discussed where we emphasize that the variable entanglement structure of the embedded fragments may vary the performance of SQD due to the lack of symmetry in the molecules, in Section 4 the results obtained are discussed and finally in Section 5 we conclude the work and mention future prospects.
@@ 2.2 Density Matrix Embedding Theory
Figure 2.1: Illustrative example of bath-orbital construction for the $H_{2}O$ molecule in the STO-3G basis. The system is fragmented so that the orbital localized on a single hydrogen atom constitutes the fragment. (a) shows an iso-surface of this localized fragment orbital; (b) shows the corresponding bath orbital obtained via DMET, which is a linear combination of the remaining localized spatial orbitals in the molecular orbital basis. Together, (a) and (b) span the impurity subspace for the H-fragment.
@@ 2.3 Sample-based Quantum Diagonalization
SQD [82] is a combination of the quantum sampling method QSCI [47] and an error-mitigation technique introduced in the work [82], S-CoRe, which “recovers” the noisy configurations. These configurations thus obtained create a sub-space, which is further expanded, and then the projection sub-space is obtained over which SCI [27] is performed to obtain the ground state energy.
@@ 3.1 DMET-SQD Workflow
    Full System Low-level Reference: The molecular co-efficient matrix $C$, and the HF determinant (in the case of single-reference systems) of the whole system are obtained using Restricted HF method (here we only encounter molecular systems with fully filled spatial orbitals).
@@
    Bath Orbital Construction: The bath orbitals are constructed with the remaining environment orbitals for the fragment $A_{y}$. Here, the occupation threshold $\epsilon_{occ}$ is set to $10^{-13}$ for discarding virtual orbitals [96].
@@
    Impurity Fock Matrix Updation: The Fock matrix $\hat{F}_{imp}$ (where $imp=A_{y}+B_{y}$) for each of the impurities is obtained, and the chemical potential $\mu_{glob}$ is subtracted from the diagonal entries corresponding to the fragment orbitals:
@@
    Impurity SCF Updation: Using $\hat{F}_{imp}^{\mathrm{new}}(\mu_{\mathrm{glob}})$ from Eq. (12), an SCF calculation is carried out in the impurity space to obtain updated molecular properties, such as the impurity coefficient matrix $C_{imp}^{\mathrm{new}}(\mu_{\mathrm{glob}})$ and the corresponding one-particle density matrix $D^{y}(\mu_{\mathrm{glob}})$. The standard convention is to initialize $\mu_{glob}^{init}=0$.
@@
    High-level Quantum Solver (SQD): Now, we use SQD, as the high-level solver to estimate the ground state energy and eigen-vector of these "impurities". The one-electron ($h_{imp}^{pq}(\mu_{glob})$) and two-electron ($h_{imp}^{pqrs}(\mu_{glob})$) Hamiltonian terms (where $pqrs\in imp$) obtained after the localization, serve as the $\hat{H}_{emb}(\mu_{glob})$, and the configurations obtained through quantum sampling (and subsequent Configuration Recovery) are used to form the subspace onto which $\hat{H}_{emb}$ is projected. In the case of calculating the reference energies, this step would be replaced by FCI, without any loss of generalization.
@@
    Fragment Energy and RDM Extraction: Using the eigen-vector obtained in the previous step, the following quantities are calculated:
@@
    Aggregation of Fragment Contributions: Now similar to the fragment $A_{y}$, the steps 3 to 7 are repeated for all the remaining $N_{frag}-1$ fragments, where $N_{frag}$ represents the number of fragments in the system. The total energy $E_{tot}(\mu_{glob})$, the number of electrons of the system $N_{tot}$, and the error in electron number $N^{ele}_{err}$ can be thus obtained by:
@@
    Chemical Potential Optimization: Next, the root of the function $N_{err}^{ele}(\mu_{glob})$ is computed to within the convergence threshold $\epsilon_{conv}$, by iteratively updating $\mu_{glob}$, and repeating the steps 3 to 8, using a root finding method (in this case we use Newton-Secant method [73]). In other words, we wish to obtain a $\mu_{glob}^{\star}$ such that:
@@ 3.2 Experimental Details
The molecules considered in this study are listed in Sec. B. The STO–3G basis set was employed to enable controlled benchmarking of the DMET-SQD framework against DMET-FCI, allowing for direct and systematic comparison in a regime where exact classical reference solutions remain tractable. In particular, the use of a minimal basis allows us to avoid additional approximations such as active space truncation in the DMET-FCI calculations, thereby ensuring a consistent and unbiased reference. Furthermore, the reduced Hilbert space complexity makes this setting well suited for studying the behavior of the embedding and sampling procedures under NISQ-era hardware constraints. Fragmentation was performed such that each atom is treated as an individual fragment for all molecules.
@@
For benchmarking, FCI is employed as it accesses the full configuration space and is both size-consistent and size-extensive. This ensures that any deviation in correlation energy arises solely from the embedding and sampling procedure, thereby providing a rigorous validation of the ability of DMET-SQD to reproduce DMET-FCI correlation within the scope of the present study.
The number of shots used for quantum sampling was $10^{4}$ and was kept consistent across all the impurities, regardless of the number of spin orbitals within them. For each impurity, 5 iterations of S-CoRe were performed. The maximum allowed samples per batch was set to $10^{5}$. The circuits were transpiled [56] with optimization set to level 3, and no additional error handling techniques were employed in these series of experiments. The convergence threshold $\epsilon_{conv}$ for $\mu_{glob}$ was set to $1.48\times 10^{-8}$, which is the default parameter in scipy.optimize.newton method. The virtual orbital threshold $\epsilon_{occ}$ was set to $10^{-13}$ for bath orbital construction. This threshold determines the inclusion of the non-negligibly entangled environment orbitals into the impurity space, thereby directly controlling the impurity size. Consequently, it influences both the amount of correlation captured and the quantum resource requirements of the corresponding SQD simulations.
@@ 4 Results and Discussion
The details pertaining to the fragmentation settings, quantum resource requirements, and the dimensions of the symmetry space and Hilbert space are summarized in Table 4.1. The resource estimation is presented in the IBM Sherbrooke superconducting hardware basis gate set {$X$, $R_{z}$, $\sqrt{X}(S_{x})$, ECRECR}. The largest quantum-sampling experiment employed 30 qubits, with a symmetry-space dimension of 25,050,025 and a Hilbert-space dimension of 1,073,741,824. The largest quantum circuit depth was 1081. These entries have been highlighted in the table to indicate the maximal values observed in our work. In contrast, the complete (unfragmented) quantum simulation of the molecules with SQD would require the following number of qubits in the STO-3G basis: $HOCN$ (32 qubits), $CH_{3}NO$ (36 qubits), $CH_{5}NO$ (40 qubits), $C_{2}H_{3}NO$ (46 qubits), $C_{2}H_{5}NO$ (50 qubits), $CH_{4}N_{2}O$ (48 qubits), $NOCl$ (38 qubits), and $HOSCN$ (50 qubits).
Figure 4.3: The localized fragment and bath orbitals for the HSOCN molecule for the fragments (a), (b), (c), (d) and (e) denoting the fragments [N], [C], [O], [H] and [S] respectively. The boxes with solid line denote the localized fragment orbitals and the boxes with dashed line denote the corresponding bath orbitals. The last three bath orbitals in subplot (e) are not considered as part of the impurity orbitals based on the value of $\epsilon_{occ}$.
The construction of bath orbitals in practical DMET implementations relies on the spectrum of the environment block of the one-particle density matrix, where only a subset of eigenvectors corresponding to nontrivial (fractional) eigenvalues are retained as entangled bath orbitals. In realistic quantum chemistry calculations, several of these eigenvalues may lie numerically very close to 0 or $1$, making it difficult to distinguish between genuinely entangled orbitals and effectively unentangled environment orbitals. Therefore, this leads to a practically feasible and efficient construction of the impurity, consisting of the localized fragment orbitals and the entangled bath orbitals with eigen-values adhering to $\epsilon_{occ}$, as discussed in the methodology, to truncate negligible contributions and define the bath space [96]. In our DMET-SQD simulations with atom-wise fragmentation, this truncation directly impacts the resulting impurity sizes. For instance, in the cases of NOCl and HOSCN, although the nominal impurity size for the [Cl] and [S] fragments would correspond to $18$ orbitals, the application of the eigenvalue cutoff reduces the effective bath to only $3$ and $6$ orbitals, respectively, yielding smaller impurity spaces consisting of $12$ and $15$ orbitals respectively, as shown in Figure 4.3 for the HOSCN molecule. This reflects that only a limited subset of orbitals significantly contributes to the fragment–environment entanglement, as determined by the chosen threshold.
From a quantum computing perspective, the choice of $\epsilon_{occ}$ plays a critical role in balancing correlation accuracy and hardware feasibility. A smaller cutoff retains more bath orbitals, thereby increasing the impurity size, which in turn raises the number of qubits and circuit depth required for SQD. This leads to a degradation in sampling quality due to increased noise accumulation on NISQ hardware, which requires more robust configuration recovery. Conversely, a larger cutoff reduces the impurity size, improving sampling efficiency and robustness, but at the cost of neglecting weaker correlations. This trade-off is particularly pronounced in low-symmetry ($C_{1}$) molecular systems, where geometric details strongly influence fragment-environment entanglement patterns. Our observations indicate that $\epsilon_{occ}$ is therefore a physically and computationally significant control knob that governs both the expressibility of the embedding and the viability of quantum sampling. Consequently, a careful, system-dependent selection of this threshold is essential to ensure that DMET-SQD calculations remain both accurate and compatible with current quantum hardware limitations. This shows that SQD within the framework of DMET is warranted further study due to the above mentioned complications.